sarahms:

Event A has a 0.3 probability of occuring and event B has a 0.4 probability of occuring. A and B are independent events. What is the probability that neither A or B occurs?

Vocaloid:

I'm not 100% sure on this but probability that neither A nor B occurs = 1 - probability that A only occurs - probability that B only occurs - probability that A and B occur probability that A and B occur = P(A) * P(B) the other probabilities are given in the problem, simply plug everything into the first equation to solve

ratnakermehta20:

\[P(A)=0.3, P(B)=0.4.\] Recall that, if A and B are independent, then, so are A' and B'. Hence, the required prob. is \[P(A'^B')=P(A')*P(B')=[1-P(A)][1-P(B)]=(0.7)(0.6)=0.42.\]